2. Compute the contracted length of an

object whose initial length 30 m and travel with a velocity 0.65c?

1
Expert's answer
2022-02-07T08:45:12-0500

Solution

initial length L=30 m

velocity v=0.65c

contracted length of an

object

"L'=L\sqrt{1-\frac{v^2}{c^2}}"

Putting all values

"L'=30\times\sqrt{1-\frac{(0.65c)^2}{c^2}}\\=30\times\sqrt{0.5775}\\=22.8m"